# Affordance-based Robot Manipulation with Flow Matching

### v3

@@ 4.2.2 Main Results
    Generation Quality: Flow matching (CNN-based, 16 steps) outperforms other baselines in terms of 2D and 3D trajectory prediction accuracy. The Transformer behavior cloning achieves marginal precision. This is expected as it is hindered by the nature of multi-modal action distribution, causing the averaging out across non-convex spaces.
  2. 2.
@@
    Stability: Fig. 7 shows an example of training and testing loss of flow matching and diffusion policy throughout the training process. We can see flow matching exhibits greater stability on both training and evaluation than diffusion policy.
  3. 3.
@@
    Inference Time: In terms of inference time, Table 2 showcases that flow matching with 16 steps achieves faster inference time compared to diffusion policy with 16 steps. We hypothesize that flow matching with linear pointwise flows (following the optimal transport theory) generates straighter flows, and thus causes faster inference. Table 2 also showcases better performances of diffusion policy when applying more inference iterations, with a trade-off of longer inference time. Contrarily, flow matching has not shown significant improvements when increasing inference steps. We hypothesize that this is because flow matching trains Continuous Normalizing Flow models, where an ordinary differential equation is solved without learning a series of discrete steps to progressively refine the generated sample. This considerably reduces the inference time for closed-loop robot manipulation, as 2-step flow matching (error: $0.888$cm, time: $8.53$ms) has achieved comparable performance as 16-step diffusion policy (error: $0.884$cm, time: $159.72$ms), but prominently lower inference time roughly by 85%.
4.2.3 Real-World Robot Experiments
@@
We deploy flow matching and other policies on real robot manipulation evaluation, as shown in Table 3. We carry out 50 replications of trials for each policy. The orientation is predefined. We use a KINOVA Gen3 arm, an Azure Kinect camera, and a 3D printed UMI gripper (Chi et al., 2024) for real-world robot experiments.
We observe that most failures occurred during the initial grasping. Even the trajectory prediction is accurate in general, a slight misalignment between the first trajectory waypoint and the object would cause grasping failures. This problem might be addressed in the future by a weighted flow matching policy, which highlights the initial waypoints by adding higher probability weights. Some out-of-distribution factors also lead to failure, including background and camera view changes. Larger data sizes and more diversity should alleviate this problem.
@@
RGB images ablations against affordances: We also investigate how the affordance would guide the flow matching policy. We trained flow matching for each task separately, taking the raw RGB images as input. It achieves a higher loss but a similar success rate compared to training all tasks in one policy with affordance guidance. Interestingly, a further comprehensive examination reinforces the argument that flow matching could handle multimodal action distribution. Fig. 8 shows an example. From the left figure, we can see that when training a separate policy of moving the towel toward the trash for sweeping, the predicted trajectory (yellow) could be detached from the ground truth (red), but still a reasonable solution that allows for a successful robot execution. With affordance guidance, the prediction is closely aligned with the truth (Fig. 8-right).
Figure 8: Ablations between using RGB images and affordances as inputs. We show examples of how visual affordance guides the flow matching policy.
4.2.4 Ablations
We further ablate various policy design choices.
@@ 4.2.4 Ablations
Network Structure: As shown in Table 2, CNN-based flow matching achieves better results than transformer-based architecture. We hypothesize that transformer might need additional hyperparameter tuning.
@@
Trajectory Horizon: We empirically test trajectory representation with 8, 16, 32 and 64 waypoints. More waypoints are not necessary, while fewer waypoints are unable to entirely encapsulate the complete long-horizon trajectories. We have found that, in general, the trajectory horizon does not wield a significant influence on flow matching performance. The performance reported in the above main results section is achieved by using 32 waypoints to represent the trajectory.
4.3 Comparisons between flow matching policy diffusion policy
To further investigate the performance of flow matching compared to diffusion policy, we benchmark the proposed methods on two more datasets: (i) Push-T adapted from (Florence et al., 2022); and (ii) Franka Kitchen proposed in Relay Policy Learning (Gupta et al., 2019).
  1. 1.
    Push-T requires pushing a T-shaped block to a fixed target with a circular end-effector. Push-T takes RGB images with proprioception of end-effector location as inputs, and end-effector actions as outputs. The experiments are carried out in a closed-loop manner. All methods are trained with $200$ demonstrations for $3,000$ epochs. Variation is added by random initial conditions for T block and end-effector.
  2. 2.
    Franka Kitchen contains 7 objects for interaction and comes with a human demonstration dataset of $566$ demonstrations, each completing 4 tasks in arbitrary order. The goal is to execute as many demonstrated tasks as possible, regardless of order, showcasing both short-horizon and long-horizon multimodality. The training takes states as inputs, and robot joint actions as outputs for $4,500$ epochs. This benchmark includes 6D closed-loop robot position and orientation actions, as well as gripper actions. Variation is added by random initial conditions for the robot and object states.
    For Push-T and Franka Kitchen benchmarks, we report results from the average of the last checkpoint with $10$ trials of replication. In addition, we also report the best perforce in the replications of the last checkpoint. The evaluation is carried out across $500$ different environment initial conditions ($5,000$ in total).
  3. 3.
    Real-world Daily Living Task, for clarity, takes affordance images as inputs, and 3D position end-effector trajectories as outputs. One-shot open-loop estimation of the entire trajectory is used in our dataset. We randomly split the dataset with $80$%-$20$% percentage of training and testing. The results reported are obtained after $1,000$ epochs of training with 3-fold cross validation to minimize the impact of any noises. Other training and testing details have been illustrated in the above sections.
Table 4 shows the hyperparameters we have used in flow matching and diffusion policy. Table 5 shows the tasks Summary.
Similar conclusions could be achieved from Table 3, as in the Main Results section: (i) Generation Quality: Flow matching outperforms diffusion policy in all benchmarks of robot manipulation experiments. (ii) Inference Time: Flow matching considerably reduces the inference time for closed-loop robot manipulation. For example, in Franka Kitchen benchmark, 2-step flow matching (success rate: $0.9750$, time: $13.18$ms) has achieved comparable performance as 16-step diffusion policy (success rate: $0.9840$, time: $157.46$ms), but prominently lower inference time roughly by 90%.
Methods Flow Matching Diffusion Policy Flow Matching Diffusion Policy Transformer Flow Matching
(Inference Step) ($2$-step) $\uparrow$ ($2$-step) $\uparrow$ ($16$-step) $\uparrow$ ($16$-step) $\uparrow$ BC $\uparrow$ separate policy
Push-T 0.8771/0.7111 0.4412/0.1872 $0.9035/0.7490$ 0.8840/0.7178
Franka Kitchen 0.9750/0.6134 0.2355/0.0527 $0.9960/0.7172$ 0.9840/0.6716
Real-world Daily Living Task $0.82$ 0.76 0.44 0.80
@@ 4.3 Comparisons between flow matching policy diffusion policy
Table 3: We present the robot performance with different checkpoint selection methods in the format of (max performance) / (average of last checkpoint with $10$ trials of replication), with each averaged across $500$ different environment initial conditions ($5000$ in total) for testing. The metric used here is success rate, except for the Push-T task which uses target area coverage.
DDPM Ta Tp ImgRes D-Net D-Para V-Enc V-Para Lr WDecay Iters Train Iters Eval
Push-T 8 16 96x96 ConditionalUnet1D 73 ResNet-18 12 1e-4 1e-6 16 2/4/8/16
Franka Kitchen 8 16 N/A ConditionalUnet1D 73 N/A N/A 1e-4 1e-6 16 2/4/8/16
Real-world Daily Living Task 32 32 224x224 ConditionalUnet1D 73 ResNet-18 12 1e-4 1e-6 16 2/4/8/16
Flow Matching Ta Tp ImgRes F-Net F-Para V-Enc V-Para Lr WDecay Iters Train Iters Eval
Push-T 8 16 96x96 ConditionalUnet1D 73 ResNet-18 12 1e-4 1e-6 N/A 2/4/8/16
Franka Kitchen 8 16 N/A ConditionalUnet1D 73 N/A N/A 1e-4 1e-6 N/A 2/4/8/16
Real-world Daily Living Task 32 32 224x224 ConditionalUnet1D 73 ResNet-18 12 1e-4 1e-6 N/A 2/4/8/16
@@
Table 4: Hyperparameters for flow matching and Diffusion Policy. Ta: action horizon. Tp: action prediction horizon. ImgRes: environment observation resolution. D-Net: diffusion network. D-Para: diffusion network number of parameters in millions. F-Net: flow matching network. F-Para: flow matching network number of parameters in millions. V-Enc: vision encoder. V-Para: vision encoder number of parameters in millions. Lr: learning rate. WDecay: weight decay. Iters Train: number of training diffusion iterations. Iters Eval: number of inference iterations. Network structures for flow matching policy are modified in the same way in diffusion policy.
@@
Real-world Daily Living Task 1 $\approx 30$ 3 8,000 N/A
Table 5: Tasks Summary. Rob: number of robots. Obj: number of objects. ActD: action dimension. PH: proficient-human demonstration. Steps: max number of rollout steps.
5 Conclusion
We have formulated a prompt tuning method for affordance learning and flow matching policy for robot manipulation. The core idea of prompt tuning is to maximally exploit the pretrained foundation model, and rapidly excavate the relevance of foundation and downstream affordance learning tasks. Our focus in this work is not to outperform state-of-the-art general robot manipulation research. Instead, we have systematically studied flow matching framework, which provides an alternative to diffusion policy. The results suggest forsaking the stochastic construction in favor of learning the probability path more directly, allowing for improved generation and faster inference. We qualitatively and quantitatively experiment on the multi-task robot manipulation scenarios to prove the ease of training and evaluation for flow matching.
The 3D estimation performs decently with our method using only RGB images as input. But we also observe the 3D estimation could be highly attuned to camera angles. We could further improve performance by incorporating depth or pointcloud information for 3D orientation modeling. We leave these for future work.

### v4

@@ IV-B2 Main Results
  • Generation Quality: Flow matching (CNN-based, 16 steps) outperforms diffusion policy and Transformer baselines in terms of 2D and 3D trajectory prediction accuracy. The Transformer behavior cloning achieves marginal precision. This is expected as it is hindered by the nature of multi-modal action distribution, causing the averaging out across non-convex spaces.
@@
  • Stability: Fig. 6 shows an example of training and testing loss of flow matching and diffusion policy with DDPM throughout the training process. We can see flow matching exhibits greater stability on both training and evaluation than diffusion policy.
@@
  • Inference Time: We have two observations here: i) Table II showcases that flow matching with 16 steps achieves faster inference time compared to diffusion policy with 16 steps. We hypothesize that flow matching with linear pointwise flows generates straighter flows than DDPM and DDIM, and thus causes faster inference. ii) More importantly, Table II also showcases that 1-step flow matching (error: $1.031$cm, time: $13.228$ms) has achieved comparable performance as 16-step DDIM (error: $2.266$cm, time: $98.680$ms), but noticeably lowered inference time roughly by 85%. We hypothesize that this is because diffusion models solve a stochastic differential equation with a series of discrete steps to progressively refine the generated sample. Contrarily, flow matching trains continuously normalize flow models, leading to no significant improvements when increasing inference steps. Thus 2-step flow matching has achieved comparable performance as 16-step diffusion policy, which considerably reduces the inference time for closed-loop robot manipulation. This is in line with some results obtained in the image generation domain. As pointed out by Stable Diffusion 3 [6], flow matching performs better with fewer inference steps than diffusion policy, especially when faster inference is required. The extensive evaluation on multiple robot benchmarks in the next section will further reinforce our argument.
  • Training Resources: DDPM training and benchmarking demand significant resources for various training and inference steps. DDIM decouples the number of denoising iterations in training and inference, thereby allowing the algorithm to be trained one time with a large training iteration and use fewer iterations for inference to speed up the process. However, flow matching still achieves faster inference than DDIM.
IV-B3 Real-World Robot Experiments
@@
We deploy flow matching, DDPM and Transformer policies on real robot manipulation for evaluation. We carry out 50 replications of trials for each baseline. We use a KINOVA Gen3 arm and an Azure Kinect camera for real-world robot experiments. Details can be found in the supplementary video.
From Table III, we can see that flow matching outperforms other baselines. We also observe that most failures occurred due to some out-of-distribution factors, including background and significant camera view changes. Larger data size and more diversity should alleviate this problem.
@@
RGB images with/without affordances: We also investigate how the affordance would guide the flow matching policy. We trained flow matching taking the raw RGB images and language tokens as input. This is similar to our proposed method but without the intermediate stage of affordance learning. It also involves some resemblance while not being entirely identical to the method in [2]. Interestingly, comprehensive examinations reinforce the argument that flow matching could handle multimodal action distribution. Fig. 7 shows one example. From the left figure, we can see that when training a policy without affordances, the predicted trajectory (yellow) of moving the towel toward the trash for sweeping could be detached from the ground truth (red), but still a reasonable solution that allows for a successful robot execution. With affordance guidance, the prediction is closely aligned with the truth (Fig. 7-right). We also observe that for applications that demand higher precision in manipulation, like grasping the toothbrush handle, affordances offer greater guidance for shaping the manipulation policy.
Figure 7: Ablations of using RGB images with/without affordances for policy training. Visual affordance guides the flow matching policy to generate a trajectory closely aligned with the truth. The policy without affordances might generate a trajectory detached from the ground truth, but potentially still a reasonable solution. This reinforces the argument that flow matching policy could handle multimodal robot action distributions.
IV-B4 Ablations
We further ablate policy design choices.
@@ IV-B4 Ablations
Network Structure: As shown in Table II, CNN-based flow matching achieves better results than transformer-based architecture. We hypothesize that transformer might need additional hyperparameter tuning.
@@
Trajectory Reprsention: We empirically test trajectory representation with 8, 16, 32 and 64 waypoints. More waypoints are not necessary, while fewer waypoints are unable to entirely encapsulate the complete long-horizon trajectories. We have found that, in general, the trajectory representation does not wield a significant influence on flow matching performance. The performance reported in the above main results section is achieved by using 32 waypoints.
Methods Flow Matching Diffusion Policy Transformer Flow Matching
(Inference Step) ($16$-step) $\uparrow$ ($16$-step) $\uparrow$ BC $\uparrow$ RGB $\uparrow$
Activities of Daily Living $0.82$ 0.76 0.44 0.74
TABLE III: Real-world robot experimental results.
Methods (16-step) Push-Ta $\uparrow$ Push-Tb $\uparrow$ Franka Kitchen $\uparrow$ Robomimic $\uparrow$
Flow Matching $0.9035/0.7519$ 0.7363/$0.6218$ $0.9960$/0.7425 $0.9360/0.7289$
DDPM 0.8840/0.7178 0.7360/0.6100 0.9840/0.6716 0.9359/0.7168
DDIM 0.8801/0.6372 $0.7490$/0.6167 0.9865/$0.7471$ 0.9334/0.7073
  • a
    sampling range: [(50, 450), (50, 450), (200, 300), (200, 300), ($-\pi,\pi$)]
  • b
    sampling range: [(50, 450), (50, 450), (100, 400), (100, 400), ($-\pi,\pi$)]
@@
TABLE IV: We present the robot evaluation performance in the format of (max performance) / (average of last checkpoint with $10$ trials of replications), with each averaged across $500$ different environment initial conditions. The metric used here is success rate, except for the Push-T task which uses target area coverage. We have used various sampling ranges (end-effector position, T-block position and orientation) for Push-T environment initialization. For Robomimic benchmark, we specifically report results on the Transport task.
IV-C Comparisons between flow matching and diffusion policy
To further investigate the performance of flow matching compared to diffusion policy, we benchmark the proposed methods on three more datasets which include closed-loop 6D robot actions and gripper actions: (i) Push-T [7], (ii) Franka Kitchen [14], and (iii) Robomimic [33].
  • Push-T requires pushing a T-shaped block to a fixed target with a circular end-effector. Push-T takes RGB images with proprioception of end-effector location as inputs, and outputs end-effector actions in a closed-loop manner. The dataset includes $200$ demonstrations.
  • Franka Kitchen contains 7 objects for interaction and comes with a human demonstration dataset of $566$ demonstrations, each completing 4 tasks in arbitrary order. The goal is to execute as many demonstrated tasks as possible, regardless of order. The training takes state-based inputs, and outputs closed-loop robot joint actions and gripper actions.
  • Robomimic consists of 5 tasks with a proficient human teleoperated demonstration dataset. We specifically focus on the transport task which includes $200$ demonstrations. The policy takes state-based inputs, and outputs closed-loop robot joint actions and gripper actions.
For each benchmark, the evaluation has been carried out across $500$ different environment initial conditions, using the last checkpoint of each policy with $10$ trials of replications. Thus, $5,000$ trials have been carried out in total per policy and benchmark. Variation is added on random initial conditions for the robot and object states. We respectively report the best and average performance in the $10$ trials of replications of the last checkpoint. All state-based tasks are trained for $4,500$ epochs, and image-based tasks for $3,000$ epochs. Table V shows the hyperparameters we have used in flow matching and diffusion policy. Table VI shows the task summary.
Similar conclusions could be achieved from Table IV and Fig. 8, as in the Main Results section:
  • Generation Quality: Flow matching outperforms DDPM in all three benchmarks. The performances of flow matching and DDIM are comparable, where flow matching performs marginally better in most cases.
  • Inference Time: Fig. 8 shows how the number of inference steps affects the performance of flow matching and diffusion policy. We can observe that diffusion policy showcases better performances when applying more inference iterations with a trade-off of longer inference time, as it requires a series of discrete steps to progressively refine the generated sample. Contrarily, flow matching has not shown significant improvements when increasing inference steps. Therefore, flow matching considerably reduces the inference time for closed-loop robot manipulation. In the Push-T benchmark in Fig. 8, 2-step flow matching (coverage: $0.8803$, time: $13.098$ms) has achieved comparable performance as 16-step diffusion policy with DDIM (coverage: $0.8801$, time: $98.268$ms), but noticeably lower inference time roughly by 86%.
Figure 8: Generation quality and inference time comparison of flow matching and diffusion policy for varying values of inference steps.
H-Param Ta Tp ObsRes F-Net F-Par V-Enc V-Par Lr WDe Iters Train (FM) Iters Train (DDPM) Iters Train (DDIM) Iters Eval
Activities of Daily Living 32 32 1x224x224 ConditionalUnet1D 72 ResNet-18 11 1e-4 1e-6 N/A 1/4/8/16 16 1/4/8/16
Push-T 8 16 1x96x96 ConditionalUnet1D 80 ResNet-18 11 1e-4 1e-6 N/A 1/4/8/16 16 1/4/8/16
Franka Kitchen 8 16 1x60 ConditionalUnet1D 66 N/A N/A 1e-4 1e-6 N/A 1/4/8/16 16 1/4/8/16
Robomimic 8 16 1x50 ConditionalUnet1D 66 N/A N/A 1e-4 1e-6 N/A 1/4/8/16 16 1/4/8/16
@@ IV-C Comparisons between flow matching and diffusion policy
TABLE V: Hyperparameters for flow matching and diffusion policy. Ta: action horizon. Tp: action prediction horizon. ObsRes: environment observation resolution. D-Net: diffusion/flow matching network. D-Par: diffusion/flow matching network number of parameters in millions. V-Enc: vision encoder. V-Par: vision encoder number of parameters in millions. Lr: learning rate. WDe: weight decay. Iters Train: number of training diffusion iterations. Iters Eval: number of inference iterations.
@@
Robomimic 2 3 20 200 700
TABLE VI: Tasks Summary. Rob: number of robots. Obj: number of objects. ActD: action dimension. PH: proficient-human demonstration. Steps: max number of rollout steps. Franka Kitchen and Robomimic involve 6D robot and gripper actions in the joint space. ADLs and Push-T focus on robot end-effector trajectories. For clarity, we further explain that for our ADLs tasks, as no closed-loop motion is considered here, we assume that the gripper closes when the first waypoint has arrived, and the low-level actions are executed between waypoints using a standard proportional-derivative (PD) controller.
V Limitations
The generation quality of flow matching and diffusion policy for robot manipulation are generally comparable. Although we see only a marginal improvement in flow matching in most cases, we would like to highlight that the focus of this work is not to outperform state-of-the-art general robot manipulation research. Instead, we have systematically studied the flow matching framework, which provides an alternative to diffusion policies for robot manipulation. We can not overlook the additional advantages of flow matching, including stable training, easy implementation, and most importantly, significantly better performance and faster inference with fewer inference steps than diffusion policy, suggesting forsaking the stochastic construction of diffusion policy in favor of learning the probability path more directly as in flow matching.
We have evaluated how varying inference steps affect the performance of flow matching and diffusion policy. The original diffusion policy research [3] adopts a large $100$ training and inference steps with DDPM in their experiments. Based on our evaluation in Fig. 8 and results from other research [4], we can observe that beyond $8$ steps, further increasing steps have only a marginal impact on the performance of diffusion policy, but with a trade-off of significantly longer inference time. We are aware of recent competitive research on one-step diffusion policy with distillation [38] and shortcut models [8]. In this paper, we primarily focus on conducting a comparative analysis of the fundamental architectures underlying flow matching and diffusion policy with DDPM and DDIM.
We have not explored applying flow matching for robot action generation in the reinforcement learning setting. Recent research [51] has modeled a policy as a return-conditional flow matching model with classifier-free guidance to eliminate many of the complexities that come with traditional offline reinforcement learning and outperformed standard imitation learning. These methods are anticipated to be seamlessly adaptable to robotic manipulation scenarios; however, comprehensive validation through practical simulations and real-world experimentation remains imperative.
VI Conclusion
We have formulated a prompt tuning method for affordance map learning and flow matching policy for robot manipulation. The core idea of prompt tuning is to maximally exploit the pretrained foundation model, and rapidly excavate the relevance of foundation and downstream affordance learning tasks. We have proposed a flow matching policy constructing paths that allow faster inference, and improved generation amongst robot behavior cloning methods. We qualitatively and quantitatively experiment on multiple robot manipulation benchmarks to prove that flow matching produces better trade-offs between computational cost and sample quality compared to prior competing diffusion-based methods.